# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ 4.4 Multi-Objective Guidance for Discrete Diffusion
In this section, we will define the multi-objective discrete diffusion problem. Then, we will introduce the concept of a Pareto optimal sequence and Monte Carlo Tree Search (MCTS). Finally, we will break down our framework of integrating MCTS to iteratively find Pareto optimal solutions using our uniform discrete diffusion model.
Pareto Optimization. When optimizing sequences for multiple objectives (e.g. affinity to multiple protein targets, stability, expression yield, etc.), there is likely no single best sequence that achieves the highest score across all objectives. Optimizing one objective often leads to sacrificing performance on another objective.
Therefore, we focus on finding a set of Pareto optimal sequences that minimize the trade-offs between objectives to achieve overall optimal performance across all objectives. Formally, Pareto-optimal sequences (or non-dominated solutions) cannot be further optimized in any single objective without sacrificing performance in another objective.
Let $\mathbf{s}(\mathbf{x})=[s_{1}(\mathbf{x}),\dots,s_{K}(\mathbf{x})]\in\mathbb{R}^{K}$ be a vector of scores that measures the performance of a sequence $\mathbf{z}$ in $K$ different objectives, with higher scores indicating better performance. A sequence $\mathbf{z}^{*}$ is said to dominate another sequence $\mathbf{x}$ (denoted as $\mathbf{x}^{*}\succ\mathbf{x}$) if and only if it satisfies the following property. For all objectives $k\in[1,K]$, the score for the $k$th objective for $\mathbf{z}^{\prime}$ is greater than or equal to the score for the $k$th objective for $\mathbf{x}$ and for at least one objective $k\textquoteright$ ’, the score for $\mathbf{x}^{*}$ is strictly greater than the score for $\mathbf{z}$.
$\small\underbrace{\mathbf{s}(\mathbf{x}^{*})\succ\mathbf{s}(\mathbf{x)}}_{``\mathbf{x}^{*}\text{ dominates }\mathbf{x}"}\;\;\;\text{iff}\;\;\;\;\underbrace{\forall k\in[1,K]\;\;s_{k}(\mathbf{x}^{*})\geq s_{k}(\mathbf{x})}_{\mathbf{x}^{*}\text{ is no worse than }\mathbf{x}\text{ in any objective}}\;\land\underbrace{\;\exists k^{\prime}\in[1,K]\;\;s_{k^{\prime}}(\mathbf{x}^{*})\geq s_{k^{\prime}}(\mathbf{x})}_{\mathbf{x}^{*}\text{ is strictly better than }\mathbf{x}\text{ in at least one objective}}$ (18)
A Pareto-optimal sequence $\mathbf{x}$ is a sequence where there does not exist another sequence $\mathbf{x}^{*}$ in the current Pareto-optimal set $\mathcal{P}^{*}$ that dominates it. Since there are trade-offs between objectives, this does not mean that $\mathbf{z}$ is dominant to all other sequences.
$\nexists\mathbf{x}^{*}\in\mathcal{P}^{*}\;\;\text{s.t.}\;\;\mathbf{s}(\mathbf{x}^{*})\succ\mathbf{s}(\mathbf{x})$ (19)
We define the Pareto frontier (PF) as the set $K$-dimensional objective score vectors corresponding to each Pareto-optimal sequence.
$\mathcal{P}^{*}=\left\{\big{(}\mathbf{x},\mathbf{s}(\mathbf{x})\big{)}\;|\;\forall\mathbf{x}^{*}\in\mathcal{V}^{L}\;\;\;\mathbf{s}(\mathbf{x})\succ\mathbf{s}(\mathbf{x}^{*})\right\}$ (20)
@@
Notation. Here, we let $\mathbf{z}_{t}$ denote the partially unmasked sequence at time $t$. $\mathbf{z}_{t}$ also corresponds to a node in the MCTS tree with a set of $M$ children nodes denoted as $\text{children}(\mathbf{z}_{t})=\{\mathbf{z}_{s,1},\dots,\mathbf{z}_{s,M}\}$ indexed from $i\in[1\dots M]$. Each child node is itself a partially unmasked sequence at time $s$ derived from sampling the MDLM reverse posterior $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$. The children nodes at each iteration of MCTS are rolled out into a set of clean sequences denoted as $\{\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,M}\}$, for each of which we compute a score vector $\mathbf{s}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$ and a rewards vector $\mathbf{r}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$, where $K$ is the number of objectives guiding the MCTS search.
@@
Let $\mathcal{P}^{*}=\{\tilde{\mathbf{x}}_{n}\}$ be the set of $|\mathcal{P}^{*}|$ Pareto-optimal sequences indexed $n\in[1\dots|\mathcal{P}^{*}|]$, which is updated at each iteration. At each node $\mathbf{z}_{t}$, we store a cumulative rewards vector $\mathbf{W}(\mathbf{z}_{t})$ and a counter for the number of times the node has been visited across all iterations $N_{\text{visit}}(\mathbf{z}_{t})$. Finally, we denote the iteration index as $i\in[1\dots N_{\text{iter}}]$, where $N_{\text{iter}}$ is the total number of search iterations.
@@
Initialization. We initialize a sequence of length $L$ with only [MASK] tokens as the root node of the MCTS tree at $t=0$ and an empty set $\mathcal{P}^{*}$ that will maintain clean sequences with Pareto-optimal score vectors. We initialize a vector of scoring functions $\mathbf{s}:\mathcal{V}^{L}\to\mathbb{R}^{K}$ that takes a clean sequence $\mathbf{x}_{s,i}\in\mathcal{V}^{L}$ generated from the partially masked sequence $\mathbf{z}_{s,i}$ and outputs a vector of real values $\mathbf{s}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$ that measures its performance in each of the $K$ objectives. We also set the hyperparameters, including the number of iterations $N_{\text{iter}}$, the number of children $M$, and the length of the token sequence $L$.
@@
At each iteration, four steps are performed to update the set of Pareto optimal solutions: traversing the tree by selecting the best child node until reaching a leaf node (selection), expanding the leaf node into $M$ distinct partially unmasked sequences (expansion), fully unmasking each child node into a clean sequence and computing multi-objective scores and rewards (rollout), and finally backpropagating the rewards to the parent nodes to guide the selection process at the next iteration (backpropagation).
@@
Selection. At each iteration, we traverse the tree starting at the root node (fully masked sequence) $\mathbf{z}_{T}$ and selecting a child node based on an edge score $\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})$ that balances child nodes that generate high reward sequences from previous iterations and unexplored unmasking actions that could lead to a larger pool of diverse sequences. From a parent node $\mathbf{z}_{t}$, we compute the selection score vector $\mathbf{U}\in\mathbf{R}^{K}$, where $K$ is the number of objectives, for each child node $\mathbf{z}_{s,i}$ by normalizing the cumulative rewards vector by the number of times the node has been visited. This ensures that we select an optimal unmasking step based on previous rollout rewards without biasing towards highly visited nodes.
$\mathbf{U}(\mathbf{z}_{s,i})=\frac{\mathbf{W}(\mathbf{z}_{s,i})}{N_{\text{visit}}(\mathbf{z}_{s,i})}$ (21)
@@
Then, we select uniformly at random from the pool of children nodes $\mathbf{z}_{s,i}\in\mathcal{P}^{*}_{\text{select}}$ whose reward vectors are non-dominant, such that there doesn’t exist another child $\mathbf{z}_{s,j}$ where the rewards across each of the $K$-objectives are equal to the reward of $\mathbf{z}_{s,i}$ and there exists a reward strictly greater than $\mathbf{z}_{s,i}$.
@@
$\mathcal{P}^{*}_{\text{select}}=\left\{\mathbf{z}_{s,i}\;|\;\nexists\mathbf{z}_{s,j}\in\text{children}(\mathbf{z}_{t})\;\;\text{s.t.}\;\;\mathbf{r}(\mathbf{z}_{s,j})\succ\mathbf{r}(\mathbf{z}_{s,i})\right\}$ (22)
@@
If the selected node is a non-leaf node, the loop repeats with the selected node $\mathbf{z}_{s,i}$ as the new parent node. Once a leaf node is reached, the loop ends and the next step executes.
@@
To ensure that the expansion step results in $M$ distinct unmasking steps, we experimented with two different batched unmasking techniques from the single partially masked sequence at a parent node. For the first method, we repeated the array corresponding to the parent node tokens over $M$ dimensions and added independently sampled Gumbel noise values $G_{s,i,j}$, where $s$ denotes the sequence in the batch, $i$ denotes the position in the sequence, and $j$ denotes the token index.
$\tilde{p}_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})$ $=\log p_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t}\big{)}+\mathbf{G}_{s}$ (23)
$\mathbf{G}_{s}$ $=-\log\left(-\log(u_{s,i,j}+\epsilon)+\epsilon\right)\;\;\;\;u_{s,i,j}\sim\text{Uniform}[0,1)$ (24)
where $\tilde{p}_{\theta}$ denotes the modified probability distribution after applying Gumbel noise and $\mathbf{G}_{s}\in\mathbb{R}^{L\times|\mathcal{V}|}$ is a matrix of independently sampled noise values.
Then, for each position $i$ in the sequence, we take the arg max index $j$ of the token with the highest probability at each position and each sequence in the batch and use it to partially unmask the tokens in $\mathbf{z}_{t}$.
@@
$\tilde{p}_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})$ $=\text{SM}\bigg{(}\text{topk}\big{\{}\log(p_{\theta}\big{(}\mathbf{z}_{s}|\mathbf{z}_{t})\big{)}+\mathbf{G}_{s}\big{\}}\bigg{)}$ (25)
Rollout. From each child node generated at time $s$, we completely unmask the sequence by greedily sampling the arg max tokens from the predicted Gumbel-max distribution $\tilde{p}_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$ at all remaining time steps from $t\to 0$ to get a set of clean sequences $\{\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,M}\}$ of SMILES tokens.
For each clean sequence $\mathbf{x}_{s,i}$, we feed it as input to the scoring functions for each of the $K$ objectives to generate the score vector $\mathbf{s}(\mathbf{x}_{s,i})=\big{[}s_{1}(\mathbf{x}_{s,i}),\dots,s_{K}(\mathbf{x}_{s,i})\big{]}\in\mathbb{R}^{K}$. Then, we use the score vector to compute a vector of rewards $\mathbf{r}(\mathbf{x}_{s,i})=\big{[}r_{1}(\mathbf{x}_{s,i}),\dots,r_{K}(\mathbf{x}_{s,i})\big{]}\in\mathbb{R}^{K}$.
The reward of a child node sequence for the $k$th objective is the fraction of the peptides in the current set of Pareto-optimal sequences $\mathcal{P}^{*}$ where the child node has a higher classifier score for the given objective. Specifically, the reward for the $i$th child node $\mathbf{z}_{s,i}$ and the resulting unmasked sequence $\mathbf{x}_{s,i}$ for the $k$th objective is computed as
$\mathbb{I}$ is an indicator function that returns 1 if the score for the $k$th objective of the child node is greater than or equal to the score of the $n$th sequence in the set of Pareto-optimal sequences. In parallel to computing the reward, we add all non-dominated children sequences to the set of Pareto optimal sequences $\mathcal{P}^{*}$ and remove all dominated sequences.
@@
In Appendix D.1, we show a proof-of-concept for a time-dependent multi-objective guidance strategy where the update to the Pareto-optimal set $\mathcal{P}^{*}$ depends on the rewards for only subset of objectives that changes depending on the current iteration, enabling the prioritization of properties with larger influence on peptide structure and function in earlier iterations and fine-tuning on additional properties in later iterations.
@@
$\mathbf{W}(\mathbf{z}_{s,i})$ $\leftarrow\mathbf{r}_{k}(\mathbf{x}_{s,i})$ (29)
@@
Then, we backtrack through the parent nodes of $\mathbf{z}_{s,i}$ up to the root node $\mathbf{z}_{T}$, adding the child reward vector to the cumulative reward vector and incrementing the number of visits for each node in the path.
$\mathbf{z}_{s,i}$ $\leftarrow\text{parent}(\mathbf{z}_{s,i})=\mathbf{z}_{t}$ (31)
@@
Penalizing Invalid Peptides. To discourage the selection process from choosing unmasking steps that result in invalid or unsynthesizable peptide SMILES, we subtract a penalty score calculated as the fraction of invalid SMILES sequences rolled out from the expanded children of a parent node determined based on our peptide SMILES validity filter described in Section 4.1. Since properties are irrelevant if the peptide SMILES is invalid, we subtract the penalty scaled by a constant $c$ from all $K$ dimensions of the cumulative reward vectors. This penalty-adjusted reward vector is backpropagated to all parent nodes to avoid paths leading to high invalid rates in the next iteration. After empirical experimentation, we determined that setting the constant $c=0.5$ leads to an increase in the fraction of valid peptides over iterations up to 100% validity.
$\mathbf{W}(\mathbf{z}_{t})$ $\leftarrow\mathbf{W}(\mathbf{z}_{t})+\sum_{i=1}^{M}\mathbf{r}(\mathbf{z}_{s,i})-c\left(\frac{N_{\text{invalid}}}{M}\right)$ (34)
Output. The output after $N_{\text{iter}}$ iterations is the set $\mathcal{P}^{*}$ of Pareto-optimal sequences across the $K$ objectives. Our strategy simultaneously guides the de-noising process towards multiple objectives while exploring the space of peptide sequences using the pre-trained unconditional MDLM without the need to approximate a conditional probability distribution or perform gradient iterations in the continuous space. Furthermore, we generate an entire set of Pareto-optimal sequences from a single pass through the model which is optimal across a total of $N_{\text{iter}}\cdot M$ total sequences sampled across all iterations.

### v2

@@ 4.4 Multi-Objective Guidance for Discrete Diffusion
In this section, we introduce the concept of Pareto dominance and non-dominance for multiple objectives and describe the Monte Carlo Tree Search (MCTS)-based algorithm for generating a set of Pareto-optimal sequences using our trained masked discrete diffusion model.
Pareto Optimization. When optimizing sequences for multiple objectives (e.g. affinity to multiple protein targets, membrane permeability, solubility, etc.), there is likely no single best sequence that achieves the highest score across all objectives. Optimizing one objective often leads to sacrificing performance on another objective.
Therefore, we focus on finding a set of Pareto optimal sequences that minimize the trade-offs between objectives to achieve overall optimal performance across all objectives. Formally, Pareto-optimal sequences (or non-dominated sequences) cannot be further optimized in any single objective without sacrificing performance in another objective.
Let $\mathbf{s}(\mathbf{x})=[s_{1}(\mathbf{x}),\dots,s_{K}(\mathbf{x})]\in\mathbb{R}^{K}$ be a vector of scores that measures the performance of a sequence $\mathbf{x}$ in $K$ different objectives, with higher scores indicating better performance. A sequence $\mathbf{x}^{*}$ is said to dominate another sequence $\mathbf{x}$ (denoted as $\mathbf{x}^{*}\succ\mathbf{x}$) if and only if it satisfies the following property. For all objectives $k\in[1\dots K]$, the score for the $k$th objective for $\mathbf{x}^{*}$ is greater than or equal to the score for the $k$th objective for $\mathbf{x}$, and for at least one objective $k\textquoteright$ ’, the score for $\mathbf{x}^{*}$ is strictly greater than the score for $\mathbf{x}$.
$\small\underbrace{\mathbf{s}(\mathbf{x}^{*})\succ\mathbf{s}(\mathbf{x)}}_{\mathbf{x}^{*}\text{ dominates }\mathbf{x}}\;\;\;\text{iff}\;\;\;\;\underbrace{\forall k\in[1,K]\;\;s_{k}(\mathbf{x}^{*})\geq s_{k}(\mathbf{x})}_{\mathbf{x}^{*}\text{ is no worse than }\mathbf{x}\text{ in any objective}}\;\land\underbrace{\;\exists k^{\prime}\in[1,K]\;\;s_{k^{\prime}}(\mathbf{x}^{*})\geq s_{k^{\prime}}(\mathbf{x})}_{\mathbf{x}^{*}\text{ is strictly better than }\mathbf{x}\text{ in at least one objective}}$ (16)
A Pareto-optimal sequence $\mathbf{x}$ is a sequence where there does not exist another sequence $\mathbf{x}^{*}$ in the current Pareto-optimal set $\mathcal{P}^{*}$ that dominates it. Since there are trade-offs between objectives, this does not mean that $\mathbf{x}$ is dominant to all other sequences.
$\underbrace{\nexists\mathbf{x}^{*}\in\mathcal{P}^{*}\;\;\text{s.t.}\;\;\mathbf{s}(\mathbf{x}^{*})\succ\mathbf{s}(\mathbf{x})}_{\mathbf{x}\text{is non-dominated}}$ (17)
We define the Pareto front as the set of non-dominated sequences $\mathbf{x}$ and their $K$-dimensional objective score vectors.
$\mathcal{P}^{*}=\left\{\big{(}\mathbf{x},\mathbf{s}(\mathbf{x})\big{)}\;|\;\nexists\mathbf{x}^{*}\in\mathcal{P}^{*}\;\;\text{s.t.}\;\;\mathbf{s}(\mathbf{x}^{*})\succ\mathbf{s}(\mathbf{x})\right\}$ (18)
@@
Notation. Let $\mathbf{z}_{t}$ denote the partially unmasked sequence at time $t$. $\mathbf{z}_{t}$ also corresponds to a node in the MCTS tree with a set of $M$ children nodes denoted as $\text{children}(\mathbf{z}_{t})=\{\mathbf{z}_{s,1},\dots,\mathbf{z}_{s,M}\}$. Each child node is itself a partially unmasked sequence at time $s$ derived from sampling the MDLM reverse posterior $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$. The children nodes at each iteration of MCTS are rolled out into a set of clean sequences denoted as $\{\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,M}\}$, for each of which we compute a score vector $\mathbf{s}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$ and a rewards vector $\mathbf{r}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$, where $K$ is the number of objectives guiding the MCTS search.
@@
Let $\mathcal{P}^{*}=\{\mathbf{x}^{*}_{n}\}$ be the set of $|\mathcal{P}^{*}|$ Pareto non-dominated sequences indexed $n\in[1\dots|\mathcal{P}^{*}|]$, which is updated at each iteration. At each node $\mathbf{z}_{t}$, we store a cumulative rewards vector $\mathbf{W}(\mathbf{z}_{t})$ and a counter for the number of times the node has been visited across all iterations $N_{\text{visit}}(\mathbf{z}_{t})$. Finally, we denote the iteration index as $i\in[1\dots N_{\text{iter}}]$, where $N_{\text{iter}}$ is the total number of search iterations.
@@
Initialization. We initialize a sequence of length $L$ with only [MASK] tokens as the root node of the MCTS tree at $t=0$ and an empty set $\mathcal{P}^{*}$ that will maintain clean sequences with Pareto-optimal score vectors. We initialize a set of scoring functions $\mathbf{s}:\mathcal{V}^{L}\to\mathbb{R}^{K}$ that takes a clean sequence $\mathbf{x}_{s,i}\in\mathcal{V}^{L}$ generated from the partially masked sequence $\mathbf{z}_{s,i}$ and outputs a vector of real values $\mathbf{s}(\mathbf{x}_{s,i})\in\mathbb{R}^{K}$ that measures its performance in each of the $K$ objectives. We also set the hyperparameters, including the number of iterations $N_{\text{iter}}$, the number of children $M$, and the length of the token sequence $L$.
@@
At each iteration, four steps are performed to update the set of Pareto optimal solutions: traversing the tree by selecting the best child node until reaching a leaf node (selection), expanding the leaf node into $M$ distinct partially unmasked sequences (expansion), fully unmasking each child node into a clean sequence and computing multi-objective score and reward vector (rollout), and finally backpropagating the total rewards to the predecessor nodes to guide the selection process at the next iteration (backpropagation).
@@
Selection. At each iteration, we traverse the tree starting at the root node (fully masked sequence) $\mathbf{z}_{T}$ and selecting a child node based on the selection score vector $\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})$ that balances child nodes that generate high reward sequences from previous iterations and unexplored unmasking actions that could lead to a larger pool of diverse sequences.
$\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})=\frac{\mathbf{W}(\mathbf{z}_{s,i})}{N_{\text{visit}}(\mathbf{z}_{s,i})}+c\cdot p_{\theta}(\mathbf{z}_{s,i}|\mathbf{z}_{t})\frac{\sqrt{N_{\text{visit}}(\mathbf{z}_{t})}}{1+N_{\text{visit}}(\mathbf{z}_{s,i})}$ (19)
The first term is the cumulative reward vector $\mathbf{W}(\mathbf{z}_{s,i})$ normalized by the number of times the node was previously visited. This guides the selection process towards the unmasking step that has resulted in fully unmasked sequences with optimal properties without biasing towards highly visited nodes. The second term is a scalar added element-wise to the normalized rewards. The scalar probability of the unmasking step based on the unconditional reverse posterior $p_{\theta}(\mathbf{z}_{s,i}|\mathbf{z}_{t})$ guides the selection towards the unmasking step with the highest probability to generate a valid peptide based on the pre-trained MDLM. When the number of times the parent node has been explored is high and the number of visits to a child node is low, the $\frac{\sqrt{N_{\text{visit}}(\mathbf{z}_{t})}}{1+N_{\text{visit}}(\mathbf{z}_{s,i})}$ term encourages exploration of the unexplored unmasking scheme given that $p_{\theta}(\mathbf{z}_{s,i}|\mathbf{z}_{t})$ is sufficiently high. However, as the number of visits to a child node increases, the impact of the second term decreases and the cumulative rewards dominate the selection score vector. $c$ is a scalar hyperparameter that determines the degree of exploration compared to exploiting high-reward nodes, which is selected to be $c=0.1$.
@@
Then, we select uniformly at random from the pool of children nodes $\mathbf{z}_{s,i}\in\mathcal{P}^{*}_{\text{select}}$ whose selection score vectors are non-dominated, such that there doesn’t exist another child $\mathbf{z}_{s,j}$ where the selection score across each of the $K$-objectives is equal to the score of $\mathbf{z}_{s,i}$ and there exists a score strictly greater than $\mathbf{z}_{s,i}$.
@@
$\mathcal{P}^{*}_{\text{select}}=\left\{\mathbf{z}_{s,i}\;|\;\nexists\mathbf{z}_{s,j}\in\text{children}(\mathbf{z}_{t})\;\;\text{s.t.}\;\;\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,j})\succ\mathbf{U}(\mathbf{z}_{t},\mathbf{z}_{s,i})\right\}$ (20)
@@
If the selected node is a non-leaf node, the loop repeats with the selected node $\mathbf{z}_{s,i}$ as the new parent node. If a fully unmasked node with $t=0$ is reached, we restart the selection process from the root node. Once a leaf node is reached, the loop ends and the next step executes.
@@
To ensure that the expansion step results in $M$ distinct unmasking steps, we experimented with two different batched unmasking techniques from the single partially masked sequence at a parent node. For the first method, we repeated the array corresponding to the parent node tokens over $M$ dimensions and added independently sampled Gumbel noise values $G^{(\ell)}_{i,j}$, where $i$ denotes the sequence in the batch, $\ell$ denotes the position in the sequence, and $j$ denotes the token index.
$\tilde{p}_{\theta,i}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t})$ $=\log p_{\theta}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t}\big{)}+\mathbf{G}^{(\ell)}_{i}$ (21)
$G^{(\ell)}_{i,j}$ $=-\log\left(-\log(u^{(\ell)}_{i,j}+\epsilon)+\epsilon\right)\;\;\;\;u^{(\ell)}_{i,j}\sim\text{Uniform}(0,1)$ (22)
where $\tilde{p}_{\theta,i}$ denotes the $i$th modified reverse transition distribution after applying Gumbel noise independently sampled for each token at each position for each sequence $i$ in the batch, where $i=[1\dots M]$. Then, for each position $\ell$ in the sequence, we sample $M$ distinct child sequences from the $M$ distinct distributions $\mathbf{z}^{(\ell)}_{s,i}\sim\tilde{p}_{\theta,i}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t}\big{)}$.
@@
$\tilde{p}_{\theta,i}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t})$ $=\text{SM}\bigg{(}\text{top}k\big{\{}\log p_{\theta}\big{(}\mathbf{z}^{(\ell)}_{s,i}|\mathbf{z}^{(\ell)}_{t}\big{)}+\mathbf{G}_{i}^{(\ell)}\big{\}}\bigg{)}$ (23)
After empirical experimentation, we found that the first method results in higher diversity across sequences whereas the second method prevents unlikely tokens. Since the reward generated by a sequence ultimately determines whether it is selected in subsequent iterations, we chose the first method to allow for greater exploration during the expansion step.
Rollout. From each child node generated at time $s$, we completely unmask the sequence by greedily sampling the argmax tokens from the predicted reverse transition distribution ${p}^{(\ell)}_{\theta,i}(\mathbf{z}_{s^{\prime}}|\mathbf{z}_{s})$ for all remaining time steps $s\to 0$ and $s^{\prime}=s-\frac{1}{T}$ to get a set of clean sequences $\{\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,M}\}$ of SMILES tokens. Then, we feed each clean sequence $\mathbf{x}_{s,i}$ as input to the scoring functions for all of the $K$ objectives to generate the score vector $\mathbf{s}(\mathbf{x}_{s,i})=\big{[}s_{1}(\mathbf{x}_{s,i}),\dots,s_{K}(\mathbf{x}_{s,i})\big{]}\in\mathbb{R}^{K}$. Then, we use the score vector to compute a vector of rewards $\mathbf{r}(\mathbf{x}_{s,i})=\big{[}r_{1}(\mathbf{x}_{s,i}),\dots,r_{K}(\mathbf{x}_{s,i})\big{]}\in\mathbb{R}^{K}$.
The reward of a child node sequence for the $k$th objective is the fraction of the sequences $\mathbf{x}^{*}_{n}$ in the current set of Pareto-optimal sequences $\mathcal{P}^{*}$ where the child node has a higher classifier score in that objective. Specifically, the reward for the $i$th child node $\mathbf{z}_{s,i}$ and the resulting unmasked sequence $\mathbf{x}_{s,i}$ for the $k$th objective is given by
$r_{k}(\mathbf{x}_{s,i})=\frac{1}{|\mathcal{P}^{*}|}\sum_{n=1}^{|\mathcal{P}^{*}|}\mathbf{1}\big{[}s_{k}(\mathbf{x}_{s,i})\geq s_{k}(\mathbf{x}^{*}_{n})\big{]}$ (24)
where $\mathbf{1}$ is an indicator function that returns 1 if the score for the $k$th objective of the $i$th child node is greater than or equal to the score of the $n$th sequence in the Pareto-optimal set $\mathcal{P}^{*}$. In parallel to computing the reward, we add all non-dominated children sequences to the set of Pareto optimal sequences $\mathcal{P}^{*}$ and remove all dominated sequences (Algorithm 6).
@@
In Appendix D.1, we show a proof-of-concept for a time-dependent multi-objective guidance strategy where the update to the Pareto-optimal set $\mathcal{P}^{*}$ depends on the rewards for only a subset of the $K$ objectives that varies depending on the current iteration, enabling the prioritization of properties with larger influence on peptide structure and function in earlier iterations and fine-tuning on additional properties in later iterations.
@@
$\mathbf{W}(\mathbf{z}_{s,i})$ $\leftarrow\mathbf{r}(\mathbf{x}_{s,i})$ (27)
@@
Then, we backtrack through the predecessor nodes of $\mathbf{z}_{s,i}$ up to the root node $\mathbf{z}_{T}$, adding the child reward vector to the cumulative reward vector and incrementing the number of visits for each node in the path. For all nodes from $\mathbf{z}_{t}=\text{parent}(\mathbf{z}_{s,i})$ to $\mathbf{z}_{t}=\mathbf{z}_{T}$ we apply the following update
@@
Penalizing Invalid Peptides. To discourage the selection process from choosing unmasking steps that result in invalid or unsynthesizable peptide SMILES, we subtract a penalty score calculated as the fraction of invalid SMILES sequences rolled out from the expanded children of a parent node determined based on our peptide SMILES validity filter. Since properties are irrelevant if the peptide SMILES is invalid, we subtract the penalty scaled by a constant $c$ element-wise from the cumulative reward vectors. This penalty-adjusted reward vector is backpropagated to all parent nodes to avoid paths leading to high invalidity rates in the next iteration. After empirical experimentation, we determined that setting the constant $c=0.5$ consistently results in 100% peptide validity rate after only $\sim$20 iterations of the MCTS search.
$\mathbf{W}(\mathbf{z}_{t})$ $\leftarrow\mathbf{W}(\mathbf{z}_{t})+\sum_{i=1}^{M}\mathbf{r}(\mathbf{z}_{s,i})-c\cdot\left(\frac{N_{\text{invalid}}}{M}\right)$ (31)
Output. The output after $N_{\text{iter}}$ iterations is the set $\mathcal{P}^{*}$ of Pareto-optimal sequences across the $K$ objectives. Our strategy simultaneously guides the unmasking process towards optimality across multiple objectives directly in the discrete state space while exploring the diverse space of peptide sequences using the trained unconditional MDLM generator. Furthermore, we generate an set of Pareto-optimal sequences from a single run through the MCTS-search algorithm which are non-dominated from the total of $N_{\text{iter}}\cdot M$ total sequences sampled across all iterations.